OpenStudy (anonymous):

The sum of the first n terms of a series is S sub n=2n^2 a. Find term 1, term 2, and term 3. b. Find S sub n - S sub (n-1)

jimthompson5910 (jim_thompson5910):

For the sequence T1, T2, T3, .... The partial sums S1, S2, S3, ... are defined as follows S1 = T1 S2 = T1 + T2 S3 = T1 + T2 + T3 etc etc So let's find S1, S2, S3 S1 = 2*(1)^2 S1 = 2*1 S1 = 2 S2 = 2*(2)^2 S2 = 2*4 S2 = 8 S3 = 2*(3)^2 S3 = 2*9 S3 = 18 So S1, S2, and S3 are S1 = 2, S2 = 8, S3 = 18 Since S1 = T1, this means that T1 = 2 Because S2 = T1 + T2, we know that S2 = T1 + T2 8 = 2 + T2 8 - 2 = T2 6 = T2 T2 = 6 and finally, since S3 = T1 + T2 + T3, we can say S3 = T1 + T2 + T3 18 = 2 + 6 + T3 18 = 8 + T3 18 - 8 = T3 10 = T3 T3 = 10 So the first three terms are 2, 6, and 10 ------------------------------------------------------- S[n] = 2n^2 S[n-1] = 2*(n-1)^2 So S[n] - S[n-1] = 2n^2 - 2(n-1)^2 S[n] - S[n-1] = 2n^2 - 2(n^2 - 2n + 1) S[n] - S[n-1] = 2n^2 - 2n^2 + 4n - 2 S[n] - S[n-1] = 4n - 2